Mark has the choice of playing two different games. When he plays Poker, the odds in favour of him winning is 2:5. When he plays Blackjack, the odds against him winning is 7:3. Which game should he play if he wants the best chances of winning? Explain your answer.
Thank you in advance for any help:)

Found 2 solutions by ikleyn, greenestamps:
Answer by ikleyn(52835)
Answer by greenestamps(13203)



Poker:
Odds in favor of winning = 2:5
Probability of winning: 2/(5+2) = 2/7

Blackjack:
Odds against winning = 7:3
Odds in favor of winning = 3:7
Probability of winning: 3/(3+7) = 3/10

3/10 is greater than 2/7; he should play blackjack to have the better chance of winning.